OpenStudy (mony01):

Products with Transposes. Use matrices to find the product of A^(T)B. A=[1] B=[1] [4] [-1]

OpenStudy (loser66):

what is A^T?

OpenStudy (mony01):

\[A ^{T}\] meaning T is transposes

OpenStudy (loser66):

Of course, I know what it is, I just ask whether you know how to get A^T from A?

zepdrix (zepdrix):

\[\Large\rm A=\left(\begin{matrix}1 \\ 4\end{matrix}\right)\]\[\Large\rm A^T=?\]

OpenStudy (mony01):

is (1 4)

zepdrix (zepdrix):

\[\Large\rm A^T=\left(\begin{matrix}1 & 4\end{matrix}\right)\]Good good good.

zepdrix (zepdrix):

\[\Large\rm \left(\begin{matrix}1 & 4\end{matrix}\right)\left(\begin{matrix}1 \\ -1\end{matrix}\right)\]Having trouble multiplying these? :o

zepdrix (zepdrix):

Might be a good idea to determine the `size` of the resulting matrix before doing any calculations.

zepdrix (zepdrix):

(1x2) multiplying a (2x1) gives us what size? :3

OpenStudy (mony01):

2x2?

zepdrix (zepdrix):

The `inner` numbers tell us whether matrix multiplication CAN BE applied. The 2's match, good, so we can multiply. The resulting matrix will be the `outer` numbers

OpenStudy (mony01):

would it be 1 and -4?

zepdrix (zepdrix):

Woops, the outer numbers are 1 and 1, yes? So our resulting matrix will be a 1x1 matrix. Yes, you've done your calculations correct, but that value should go into our single slot. 1(1)+4(-1)

OpenStudy (mony01):

-3?

zepdrix (zepdrix):

\[\Large\rm \left[\begin{matrix}1 & 4\end{matrix}\right]\left[\begin{matrix}1 \\ -1\end{matrix}\right]=[-3]\]Yay good job \c:/

OpenStudy (mony01):

thank you for your help

OpenStudy (mony01):

hey @zepdrix do you know when it will not be compatible?

zepdrix (zepdrix):